OpenStudy (anonymous):

Karen has scored 28, 27, and 25 points in her three basketball games so far. How many points does she need to score in her next game so that her average (mean) is 27 points per game?

OpenStudy (anonymous):

Average = total 4 games/ 4 @saracanada Can you plug the numbers in?

OpenStudy (anonymous):

i don't know??

OpenStudy (anonymous):

What's the average?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Yes, then what is the 3 scores?

OpenStudy (anonymous):

*are

OpenStudy (anonymous):

28, 27, and 25

OpenStudy (anonymous):

The fourth score is x!

OpenStudy (anonymous):

What's the total of 3 scores?

OpenStudy (anonymous):

80

OpenStudy (anonymous):

Now plug into the formula: Average = total 4 games/ 4

OpenStudy (anonymous):

20

OpenStudy (anonymous):

or 4 * 27 = 80 + x Can you find x?

OpenStudy (anonymous):

108

OpenStudy (anonymous):

Why don't you write into the equation, instead of giving one by one random number?

OpenStudy (anonymous):

4*27=108 108+80=1.35

OpenStudy (anonymous):

Because you don't learn to plug into the equation, your result is wrong: 4 * 27 = 80 + x 108 - 80 = x

OpenStudy (anonymous):

108-80=28

OpenStudy (anonymous):

Your goal is find x ( the fourth score), then x needs to be in the equation: 4 Average = total 4 games 4 * 27 = 80 + x 108 - 80 = x Thus x = 28 Learn to write algebra properly, will you :/

OpenStudy (anonymous):

thank u i will study hard

OpenStudy (anonymous):

It's just 2 lines of plugging to the formula :)